### v2

# An Inversion-based Measure of Memorization for Diffusion Models

Abstract
The past few years have witnessed substantial advances in image generation powered by diffusion models. However, it was shown that diffusion models are vulnerable to training data memorization, raising concerns regarding copyright infringement and privacy invasion. This study delves into a rigorous analysis of memorization in diffusion models. We introduce an inversion-based measure of memorization, InvMM, which searches for a sensitive latent noise distribution accounting for the replication of an image. For accurate estimation of the memorization score, we propose an adaptive algorithm that balances the normality and sensitivity of the inverted distribution. Comprehensive experiments, conducted on both unconditional and text-guided diffusion models, demonstrate that InvMM is capable of detecting heavily memorized images and elucidating the effect of various factors on memorization. Additionally, we discuss how memorization differs from membership. In practice, InvMM serves as a useful tool for model developers to reliably assess the risk of memorization, thereby contributing to the enhancement of trustworthiness and privacy-preserving capabilities of diffusion models.
1 Introduction
[…]
However, large-scale data sets used to train the prevalent DMs, e.g., the open-source image-caption dataset LAION [45], are widely acknowledged to contain content that will raise concerns regarding copyright infringement, privacy invasion, etc. For example, recent reports revealed that LAION could refer to photographers’ work without authorization [18] and contain private medical photographs [2]. With the uncurated data for training, diffusion models are likely to generate content that infringes the creators’ rights or exposes private information [9, 46]. In this work, we focus on memorization in DMs, one of the most critical issues of training data misuse.
Figure 1: Schematic illustration of diffusion model inversion, optionally with conditions.
Memorization usually manifests itself at sampling time, when a model generates images closely resembling training images. It can turn into a problem both technically and practically. Technically, memorization indicates failure of generalization to some degree because a probabilistic generative model is supposed to produce novel and diversified images rather than simply replicate those in the training set. Practically, memorization might also raise ethical and legal concerns. First, even the model developers are authorized to train their model with protected images, the image owners will never expect their works to be exposed to arbitrary users due to memorization as this would cause indisciplinable dissemination, violating the contextual integrity [36]. Second, in past years, text-to-image models have been facing lawsuits for generating derivative images that mimic the style of artists. Compared to derivative generations whose legality is still in pending [44], exact replication of copyrighted images is undisputedly intolerable. Finally, a series of works [24, 37] have proposed to use synthetic data in place of real data to prevent sharing of private information. For this goal, potential memorization should also be carefully circumvented.
Recent studies [9, 50, 51, 57] confirm the existence of memorization in DMs in a two-stage way. They first randomly draw samples from the model and then detect risky generations of low diversity [9], search for nearest neighbors in the training set [50, 51] or identify abnormal prediction errors [57]. However, conducting a practical analysis of memorization with these approaches still faces challenges. Transforming the memorization of an image to the presence or absence in samples from the model is not a reliable and efficient solution. While certain images do not appear in a set of samples, it does not demonstrate that they are not highly memorized [54]. For DMs with large latent space, adequate samples should be tested to ensure as complete coverage of the latent space as possible, which usually requires up to millions of testings [57, 9]. Furthermore, things will get much more complicated in conditional DMs, e.g, text-to-image generation. The additional conditions expand the input space of DMs, making the memorization of an image depend on the specific condition. Recent studies use training captions to induce the model to emit corresponding highly memorized training images. However, it still remains unknown that whether the other images not triggered by their coupled captions are also to high extent of memorization.
To address these challenges, we propose an Inversion-based Measure of Memorization (InvMM [^1] The source code can be found at https://github.com/Maryeon/InvMM.) for DMs, an continuous instance-level measure that takes an image as input, outputs a scaler quantifying its level of memorization. As shown in Fig. 1, the inversion process locates a sensitive latent noise distribution and optionally another condition distribution for the target image, such that generating with noise and condition drawn from the inverted distributions will produce samples exactly the same as the target image. InvMM is then defined as the minimum KL divergence to the standard Gaussian distribution and estimated through an adaptive optimization algorithm. Compared to existing studies, InvMM is more reliable and efficient by explicitly searching for all potential sensitive inputs and avoiding to collect any conditions (e.g., prompts). It enables us to discover highly memorized images, claim the safety against memorization for normal images, determine the best training setting, etc. In summary, we make the following contributions in this paper:
  • We propose a novel measure of memorization for DMs, that is more reliable and efficient.
  • For accurate estimation of the measure, we provide an adaptive algorithm that alleviates overestimation and underestimation.
  • We validate the effectiveness of our measure through comprehensive experiments across four datasets, encompassing both unconditional and text-guided DMs.
2 Related work
2.1 Memorization of generative models
Memorization has raised wide attention on various generative models, including GAN [17], VAE [29], language models [7] and DM [49]. There have been studies on training algorithms [43, 4, 22] and evaluation protocols [19] to improve the generalization ability of generative models that would potentially circumvent memorization. Empirical and theoretical studies also find that dataset complexity and size [15, 50], training time [53], training data duplication [10, 59] and model capacity [5, 10] are influential factors of memorization. Several works [53, 58] focus on the latent space (noise space) to analyze memorization. Nagarajan et al. [53] demonstrate that when trained on sufficient random noise vectors, a generator is guaranteed to memorize training data. Webster et al. [58] measure overfitting of GANs via the reconstruction error at the optimal noise vector. We take a similar perspective on the latent space to analyze memorization while aim to reach an instance-level measure of memorization rather than reveal overall properties of models. In addition, Van der Burg et al. [54] also propose an instance-level measure of memorization as the changed likelihood of one sample when removing it from the training set. However, as discussed by them, the measure is not necessarily comparable between datasets, which is important to understand the true level of memorization.
Specifically for diffusion models, recent research [9, 50, 51] demonstrates that memorization occurs in text-to-image diffusion models. Some heavily memorized images are discovered from numerous generated samples by membership inference [9] or searching for the most similar training images with image matching models [50]. Subsequently, Webster [57] provides more efficient attacks to extract training images from text-to-image models. Wen et al. [61] focus on the detection of abnormal prompts that will trigger generation of training images. In response to the disclosed threats, Vyas et al. [55] propose a provable copyright protection method to prevent replication of sensitive training images. The model is trained to align with a safe model that does not take sensitive data for training. Chen et al. [12] prevent training data replication by a modified sampling procedure that dynamically twists the sampling direction far away from any training points. Our method complements existing studies as a effective way to quantify memorization for any sample (not necessarily training samples).
2.2 Inversion of diffusion models
Inversion techniques in diffusion models are widely studied for image editing [16, 35, 63], copyright protection [62], generated content detection [56, 11], etc. Through inversion, the object, style and concept contained in the source images can be compressed in latent noises or input token embeddings (for text-to-image generation). Then the inverted latent noises or input token embeddings are utilized to generate same/similar images that preserve the desired content.
We leverage inversion techniques to analyze memorization in diffusion models. In contrast, we invert a sample to a distribution consisting of all feasible latent vectors rather than to a single vector [16, 35, 62, 56, 11]. We also adopt more efficient training objective for inversion, without the need to iteratively invert a specific sampling procedure [35, 62, 56, 11].
3 Background
Basically, DMs [49, 22] are likelihood-based generative models that learn to maximize the likelihood $p_{\bm{\theta}}(\bm{x}_{0})$ of observed data $\bm{x}_{0}$. Directly calculating $p_{\bm{\theta}}(\bm{x}_{0})$ is intractable for continuous data such as image. By contrasting to a sequence of latent variables $\bm{x}_{1:T}$, the log-likelihood $\log p_{\bm{\theta}}(\bm{x}_{0})$ can be maximized indirectly by maximizing its variational lower bound [28, 33]:
[…]
$\geq\mathbb{E}_{q(\bm{x}_{1:T}|\bm{x}_{0})}\left[\log\frac{p_{\bm{\theta}}(\bm{x}_{0:T})}{q(\bm{x}_{1:T}|\bm{x}_{0})}\right]$ (2)
Assume both transitions $p_{\bm{\theta}}(\bm{x}_{0:T})$ and $q(\bm{x}_{1:T}|\bm{x}_{0})$ are Markov chains, i.e., $q(\bm{x}_{1:T}|\bm{x}_{0})=\prod_{t=1}^{T}q(\bm{x}_{t}|\bm{x}_{t-1})$, $p_{\bm{\theta}}(\bm{x}_{0:T})=p(\bm{x}_{T})\prod_{t=1}^{T}p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$, the variational lower bound in Eq. 2 expands as:
$\log p_{\bm{\theta}}(\bm{x}_{0})\geq\mathbb{E}_{q(\bm{x}_{1}|\bm{x}_{0})}\left[\log p_{\bm{\theta}}(\bm{x}_{0}|\bm{x}_{1})\right]$
$-\sum_{t=2}^{T}\mathbb{E}_{q(\bm{x}_{t}|\bm{x}_{0})}\left[D_{\mathrm{KL}}(q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})\;\|\;p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t}))\right]$ (3)
$-D_{\mathrm{KL}}(q(\bm{x}_{T}|\bm{x}_{0})\;\|\;p(\bm{x}_{T}))$ (4)
where $p(\bm{x}_{T})$ is generally assumed to be a simple distribution such as standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$.
The transition $q(\bm{x}_{1:T}|\bm{x}_{0})$ is not learned but replaced by a manually crafted non-parametric diffusion process. The transition $p_{\bm{\theta}}(\bm{x}_{0:T})$ is the corresponding reverse (or generation, sampling) process. The diffusion process gradually adds Gaussian noises to the input $\bm{x}_{0}$ according to a weight schedule $\alpha_{1:T}$:
$\bm{x}_{t}=\sqrt{\alpha_{t}}\bm{x}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}_{t-1},\bm{\epsilon}_{t-1}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (5)
$\bm{x}_{t}=\sqrt{\bar{\alpha}_{t}}\bm{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}_{0},\bm{\epsilon}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (6)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$ decreases over time and finally to almost zero at the last step $T$, such that $q(\bm{x}_{T}|\bm{x}_{0})$ approaches to a standard Gaussian and thus Eq. 4 approximately equals to 0.
If the diffusion process is divided into sufficient steps, each reverse transition $p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t})$ can also be approximated by a Gaussian transformation with learnable parameters that is trained to match the corresponding posterior diffusion transition $q(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})$. Omitting further details [33], the following simplified objective [22] as an estimate of the variational lower bound is commonly used for training:
$l_{de}=\mathbb{E}_{\bm{\epsilon}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),t\sim\mathcal{U}\{1,T\}}\left[\lVert\bm{\epsilon}_{0}-\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\rVert_{2}^{2}\right]$ (7)
where $\bm{\epsilon}_{\bm{\theta}}$ is a neural network that takes the noisy data $\bm{x}_{t}$ and timestep $t$ as input and predicts the added noise $\bm{\epsilon}_{0}$. After training, the vanilla sampling procedure starts with a random Gaussian noise $\bm{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and removes the predicted noise stepwise. However, it is extremely slow to generate an image as it must invoke the network $\bm{\epsilon}_{\bm{\theta}}$ for $T$ times (typically 1000 steps). To mitigate the problem, a variety of efficient sampling algorithms are proposed, such as DDIM sampler [52] and PLMS sampler [32].
4 Inversion-based measure
Given an image $\bm{x}_{0}$, we aim to measure to what extent $\bm{x}_{0}$ is memorized by a diffusion model (DM). We propose to invert the image into a latent noise distribution, such as noise vectors drawn from it will lead to replication of the image $\bm{x}_{0}$. Our memorization measure is established in this inversion process. In this section, we introduce a sensitivity definition of the inverted noise distribution (Sec. 4.1), elaborate the memorization measure (Sec. 4.2) and give an adaptive algorithm for its calulation (Sec. 4.3). Ultimately, we present an example to measure memorization in conditional DM, specifically text-to-image DM (Sec. 4.4).
4.1 Sensitivity of latent distribution
The diversity of DM generated content lies in the randomness of latent noise vectors. Starting from different noise vectors, the generated content generally vary. In the context of memorization, this motivates us to search for sensitive noise vectors that will replicate the target image. We then figure out “how many” sensitive noise vectors there are and adopt it as a meaningful measure.
Considering the fact that the common latent noise is sampled from a standard Gaussian distribution, we correspondingly assume the sensitive noise vectors follow a distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ with learnable parameters $\bm{\varphi}$. The sensitivity function $\mathcal{S}(q_{\bm{\varphi}})$ of a noise distribution can be defined as the probability that generated image $G(\bm{\epsilon}_{0})$ obtained from random noise $\bm{\epsilon}_{0}\sim q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ replicates the target image $\bm{x}_{0}$:
[…]
where $\mathbf{1}_{v}$ is an indicator function that equals 1 if v is true and 0 otherwise, $d(\cdot,\cdot)$ is a distance metric between two images and $\beta$ is an appropriate threshold.
The choice of the distance metric depends on specific applications. In applications that accept derivative generation, a strict metric that only detects eidetic generation can be utilized. While under some circumstances even similar generation should be circumvented, a looser metric is needed. In this work, we consider exact replication that the generated image and target image are exactly the same, following previous work [9]. This consideration to some extent reduces the ambiguity of determining the similarity of two images. Following the investigation of Somepalli et al. [50], we implement the metric as the Euclidean distance between embeddings extracted by the pre-trained copy detection model SSCD [38].
4.2 Quantify memorization
Based on the definition of sensitivity, we propose an Inversion-based Measure of Memorization (InvMM), as the best achievable normality of $q_{\bm{\varphi}}(\bm{\epsilon}|\bm{x})$ under the condition that $\mathcal{S}(q_{\bm{\varphi}})=1$:
$\min_{\bm{\varphi}}D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x})\;\|\;p(\bm{\epsilon}_{0}))$
$\mathrm{s.t.}~{}\mathcal{S}(q_{\bm{\varphi}})=1$ (9)
where $p(\bm{\epsilon}_{0})$ is the prior distribution used in DM, which is often selected to be standard Gaussian. The normality of inverted noise distribution is the KL divergence to the prior.
(a) Optimal
(b) Overestimate
(c) Underestimate
Figure 2: Illustration of our memorization measure. The gray blocks indicate the distribution of sensitive latent noise vectors.
Explaining InvMM. Partitioning the normality gives us a more intuitional understanding of the measure. Notice that
$D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})\;\|\;p(\bm{\epsilon}_{0}))=$
$\mathrm{H}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0}),p(\bm{\epsilon}_{0}))-\mathrm{H}(q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0}))$ (10)
[…]
InvMM requires to achieve the best trade-off between improving normality and meeting the sensitivity condition. Different images to varied extent of memorization may reach different trade-off. Figure 2 illustrates the optimal solution compared to suboptimal choices. Overestimate of memorization produces a distribution with good normality but low sensitivity (Fig. 2(b)). Conservative inversion of the latent distribution perfectly meets the sensitivity condition but underestimates memorization (Fig. 2(c)).
4.3 Adaptive algorithm
We then describe how to solve the inversion problem and introduce an adaptive algorithm to calculate the proposed measure.
Transforming the inversion problem. It is intractable to directly calculate the sensitivity function. For diffusion models, it can be achieved by minimizing the denoising error $l_{de}$ (see Eq. 7) as a proxy. As can be expected, when $l_{de}$ decreases to a certain level, $\mathcal{S}(q_{\bm{\varphi}})$ will approach 1.
Moreover, denote the KL divergence in Eq. 9 by $l_{kl}$, the optimization problem can be solved by Lagrangian method:
$\min_{\bm{\varphi},\lambda}l_{de}(\bm{x}_{0};\bm{\varphi})+\lambda l_{kl}(\bm{x}_{0};\bm{\varphi}),\lambda>0$ (11)
where $\lambda>0$ is a multiplier. Note that we attach the multiplier to $l_{kl}$ to ensure adequate optimization of the denoising error $l_{de}$.
Different from the standard $\bm{\epsilon}_{0}$-prediction denoising error, we adopt $\bm{x}_{0}$-prediction denoising error, which shows more stable inversion efficacy in practice:
$l_{de}$ $=\mathbb{E}_{q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0}),t}\left[\lVert\bm{x}_{0}-\bm{x}_{\bm{\theta}}(\bm{x}_{t},t)\rVert_{2}^{2}\right]$
$=\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}\mathbb{E}_{q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0}),t}\left[\lVert\bm{\epsilon}_{0}-\bm{\epsilon}_{\bm{\theta}}(\bm{x}_{t},t)\rVert_{2}^{2}\right]$ (12)
where $\bm{x}_{t}$ is obtained as in Eq. 6 with added noise drawn from $q_{\bm{\varphi}}$ rather than standard Gaussian.
For the noise distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$, we assume it to be a multivariate Gaussian $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ with learnable mean $\bm{\mu}$ and diagonal variance $\bm{\sigma}^{2}$, i.e., $\bm{\varphi}=(\bm{\mu},\bm{\sigma}^{2})$. Then InvMM has a closed-form expression $\frac{1}{2}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$. Although the true distribution of sensitive latent noise might not be fully captured by a Gaussian distribution, we choose this modeling for simplicity and find it effective in experiments.
Algorithm 1 InvMM: adaptive weight tuning algorithm.
0:  Noise prediction network $\bm{\epsilon}_{\bm{\theta}}$, target image $\bm{x}_{0}$, optimization iteration $S$, observation cycle $C$, weight increment $\delta$, threshold $\xi$, learning rate $\gamma$
1:  Initialize $\bm{\varphi}\leftarrow(\bm{\mu}\leftarrow\mathbf{0},\bm{\sigma}\leftarrow\mathbf{I}),\lambda\leftarrow 1,l_{de}^{p}\leftarrow+\infty,\mathrm{success}\leftarrow\mathrm{FALSE}$
2:  for $i\leftarrow 1$ to $S$ do
3:     $\mathcal{L}\leftarrow l_{de}(\bm{x}_{0};\bm{\varphi})+\lambda l_{kl}(\bm{x}_{0};\bm{\varphi})$
4:     $\bm{\varphi}\leftarrow\bm{\varphi}-\gamma\nabla_{\bm{\varphi}}\mathcal{L}$
5:     if $i>1~{}\textbf{and}~{}i~{}\%~{}C=0$ then
6:        if $l_{de}-l_{de}^{q}<\xi$ then
7:           $\lambda\leftarrow\lambda~{}/~{}2$
8:        else
9:           $\lambda\leftarrow\lambda+\delta$
10:        $l_{de}^{q}\leftarrow l_{de}$
11:        if $\mathcal{S}(q_{\bm{\varphi}})=1$ then
[…]
13:           break
14:     else
15:        $\lambda\leftarrow\lambda+\delta$
16:  if success then
17:     return  $\frac{1}{2}\sum_{i}(\bm{\mu}^{2}_{i}+\bm{\sigma}^{2}_{i}-\log\bm{\sigma}^{2}_{i}-1)$
18:  else
19:     return  +$\infty$+ ∞
Adaptive weight tuning. To achieve the best trade-off, the weight $\lambda$ should be adjusted appropriately for different samples. However, it cannot be optimized via gradient decent together with $\bm{\varphi}$. As the gradient w.r.t. $\lambda$ is $l_{kl}$, which is always nonnegative, $\lambda$ will gradually decrease to 0 and $l_{kl}$ will be totally ignored. To solve the problem, we propose an adaptive weight tuning algorithm that allows dynamic change of $\lambda$ during optimization, based on an additive increase/multiplicative decrease scheme similar to congestion control of TCP protocol, as in Algorithm 1, line 6-9.
Specifically, we monitor the denoising error $l_{de}$ every $C$ steps during the optimization process and divide the weight $\lambda$ by half if no significant improvement of denoising error is observed in the cycle. Otherwise, $\lambda$ slowly increases by a small increment $\delta$ each step. In this way, $\lambda$ is adaptively adjusted for different samples to balance the normality and denoising error. When the normality is too strong to further reduce the denoising error, thus to meet the sensitivity condition, it is immediately down-weighted. Otherwise, we gradually strengthen it for better normality. We also perform early stop (line 11-13) when the sensitivity condition is reached to avoid unnecessary normality degradation (underestimate). The sensitivity condition is tested via the Monte Carlo method. Random latent vectors are drawn and used to generate samples. If the distance between any of them and the target image is lower than the threshold $\beta$, then the inversion is successful. There are cases that only partial samples replicate that target image in the end. We still consider such cases as insignificant memorization and flag them as unsuccessful.
4.4 Text-to-image DM
For text-to-image DM, we model the prompt part as a distribution as well and optimize it simultaneously with the noise distribution. Suppose each input prompt is tokenized into maximum $N$ tokens $\bm{\omega}=[\bm{\omega}_{1},\bm{\omega}_{2},...,\bm{\omega}_{N}]$. Motivated by Guo et al. [20], we learn a $|\mathcal{V}|$-dimensional categorical distribution over the vocabulary $\mathcal{V}$ for each token $\bm{\omega}_{i}$. Let token $\bm{\omega}_{i}$ follows a categorical distribution with probabilities $\bm{\pi}_{i,j}$, where $\sum_{j=1}^{|\mathcal{V}|}\bm{\pi}_{i,j}=1$. $\bm{\pi}_{i,j}$ can be parameterized and optimized.
Because sampling from categorical distribution is non-differentiable, we adopt the Gumbel-Softmax reparameterization [23] to obtain smoothed sample $\tilde{\bm{\omega}}_{i}$. Given the smoothed token $\tilde{\bm{\omega}}_{i}$, it can be converted to token embedding as a linear combination of vocabulary embeddings $e(\tilde{\bm{\omega}}_{i})=\sum_{j=1}^{|\mathcal{V}|}\tilde{\bm{\omega}}_{i,j}e(j)$, where $e(j)$ is the pretrained token embedding of the $j$-th token. A proper choice of the prompt distribution is critical for inversion on text-to-image DM. We provide a comprehensive analysis in the supplementary material.
5 Experiment
In this section, we conduct a series of experiments from different perspectives to demonstrate the validity of InvMM. First, we conduct controlled studies to verify the effectiveness of the adaptive algorithm (Sec. 5.2). Based on this, we show that InvMM can accurately detect highly memorized images (Sec. 5.3) and elucidate the influence of various factors on memorization (Sec. 5.4). Finally, we discuss how memorization differs from membership (Sec. 5.5).
5.1 Experiment setup
Dataset. We use four datasets for evaluation, i.e., CIFAR-10 [30], CelebAHQ [25], FFHQ [26] and LAION-Aesthetics V2 [45]. For subsequent experiments, we construct four randomly sampled subsets of CelebAHQ and FFHQ, CelebAHQ-250, CelebAHQ-2.5k, FFHQ-600 and FFHQ-6k. We ensure that the larger subset includes the smaller subset. The LAION dataset is extremely large. To keep computational costs manageable, we collect a subset of images. We first include 76 heavily memorized images discovered by Webster [57]. We call this set of images the confirmed set. Furthermore, we collected another 50 images with high duplication rates according to the statistics from Webster et al [59]. This set of suspects, named as the suspicious set, are assured to be disjoint with the confirmed set. Images in the confirmed set can be triggered by their training captions while those in the suspicous set cannot. Finally, another 50 randomly sampled images are used for comparison, referred to as the normal set.
Diffusion Models. Following the implementations of DDPM [22] and Latent Diffusion Model (LDM) [41], we train DDPM on CIFAR-10 and LDM on the subsets of CelebAHQ and FFHQ. Pre-trained LDMs on the whole CelebAHQ and FFHQ are also used. For text-to-image generation, we utilize Stable Diffusion (SD) v1.4 and v2.1 trained on LAION.
Inversion. The experiment setting and additional results, including prompt inversion for SD, are deferred to the supplementary material.
5.2 Adaptive algorithm
Figure 3: The influence of weight $\lambda$ on measuring memorization.
(a) CIFAR-10 (DDPM)
(b) LAION (SD)
(c) CelebAHQ (LDM)
(d) FFHQ (LDM)
Figure 4: Random samples generated using latent noise vectors drawn from inverted distributions. The original training images are in the first column within each block.
The core design of Algorithm 1 is the dynamic reweighting of the two loss items. To verify its effectiveness, we compare to fixed weights $\lambda=1$ and $\lambda=0$. $\lambda=1$ equals DM’s original training objective (Eqs. 3 and 4) and $\lambda=0$ totally ignores the normality constraint.
A comparison of the resulting memorization score under different settings can be observed in Fig. 3. Fixing $\lambda$ to 1 cannot successfully invert most images: 92.4% and 99.7% images of CelebAHQ and FFHQ are assigned infinite memorization scores. In contrast, both $\lambda=0$ and adaptive reweighting successfully invert all images. However, Fixing $\lambda$ to 0 leads to larger resulting memorization scores, which indicates underestimation of the extent of memorization.
Figure 4 illustrates the generation results with the learned distributions. With the adaptive algorithm, we are able to successfully perform inversion on various datasets and DMs. The random generated samples closely resembles the training images. The results support the premise of InvMM being a meaningful measure.
5.3 Detection
Metric DDPM (CIFAR-10) SD v1.4 (LAION)
AUC TPR@1%FPR AUC TPR@1%FPR
$\bm{\epsilon}_{0}$-loss 0.981 0.414 1.000 1.000
$\bm{x}_{0}$-loss 0.180 0.000 0.992 0.961
$M^{\mathrm{LOO}}$ [54] 0.418 $0$
Wen [61] - - 1.000 1.000
InvMM* - - 1.000 1.000
InvMM 0.997 0.899 0.994 0.842
Table 1: Detection performance on DDPM and SD v1.4. *: using training captions, fixed during inversion.
For InvMM to be a correct measure of memorization, it should be capable of distinguishing highly memorized images from normal ones. To construct a reliable test set, we conduct a nearest neighbor test following previous studies [50, 58, 9] to discover highly memorized images from DM’s training set. In this sense, highly memorized images consist of training images that can appear at sampling time. Specifically, we consider DDPM trained on CIFAR-10, LDMs trained on CelebAHQ and FFHQ, and SD trained on LAION. With their official implementations, only DDPM and Stable Diffusion v1.4 present obvious replications, which are thus used for evaluation. Finally, for DDPM, the 99 discovered training images appearing in 1M random samples are used as positive samples and another 1k different training images are included as negative samples; for SD v1.4, we deem images in the confirmed set as positive and those in the normal set as negative.
Table 1 presents the detection performance assessed via AUC and TPR@1%FPR. For the unconditional DDPM on CIFAR-10, training loss metrics (the first 2 rows) show decent results and the original $\bm{\epsilon}_{0}$-loss works much better than the $\bm{x}_{0}$-loss. The differential metric $M^{\mathrm{LOO}}$ is less effective than the training loss and InvMM. The results indicate that training loss based metrics (including $M^{\mathrm{LOO}}$) are not reliable to signify memorization, as they are not fundamentally commensurable between samples. InvMM measures memorization of each sample by contrasting their sensitive noise distribution to a common standard Gaussian, thus having the same scale.
For SD v1.4, we compare to training loss metrics and the metric proposed by Wen et al. [61]. As the compared methods all require access to image captions, we implement the counterpart for InvMM as well. With access to captions, all the metrics achieve perfect detection performance. However, this does not mean that the compared methods can also be utilized as memorization measures. On the one hand, training captions might not be available in practice, e.g., for non-training images. On the other hand, exploiting training captions for detection is actually to predict whether the captions will replicate the training images, rather than measure the memorization. In contrast, that our measure calculated with prompt optimized simultaneously shows a degree of performance drop, implies that images considered as less memorized according to the behavior of their training captions might be actually memorized more, exposed by potential prompts. In this sense, our measure provides a worst-case estimate of memorization for conditional DMs.
5.4 Influence factors
Both theoretical analysis and empirical investigation point out that dataset complexity [15, 51], dataset size [15, 51], cardinality of latent vector set [53] and data duplication [59] are important factors for memorization. We therefore verify the correctness of our measure by utilizing it to describe tendency of memorization in the change of these factors.
Dataset complexity, size and training epoch. First of all, we simultaneously validate dataset complexity, size and cardinality of latent vector set on CelebAHQ and FFHQ. Following Feng et al. [15], we measure dataset complexity with Intrinsic Dimensionality (ID) [31]. The cardinality of latent vector set can be controlled by training epochs (if in each iteration a fresh random latent vector is picked). However, we argue that previous works [15, 51] obfuscate the effect of dataset size and training epochs, as models are trained for different epochs (probably more) on smaller subsets to the full dataset. Consequently, on smaller subsets, each training image would see different number of latent vectors. Therefore, in our experiment, we decouple these two factors by training for the same epochs as on the full dataset when studying the influence of dataset size.
Figure 5: Dataset complexity & size vs. memorization. All the models are trained for 943 epochs on CelebAHQ and 446 epochs on FFHQ.
Figure 6: Training epoch vs. memorization. Models are trained for multiples of the default epoch (CelebAHQ: 943, FFHQ: 446).
As can be observed in Fig. 5, dataset complexity and size seem not present high correlation with memorization, as indicated by our measure. However, when trained for longer time, the memorization scores gradually decrease, which indicates a large extent of memorization. Therefore, dataset complexity and size are not independent determining factors, and in conclusion, our measure accurately reflects the impact of these three factors on memorization.
Figure 7: Duplication vs. memorization on CelebAHQ-2.5k and FFHQ-6k. CelebAHQ-250 and FFHQ-600 subsets are duplicated 5, 10, 15 and 20 times respectively. Outliers are hidden for clarity.
Figure 8: Deduplication vs. memorization.
Data duplication. We then validate the effect of data duplication in DMs’ training set through our measure. We train a series of LDMs on CelebAHQ-2.5k and FFHQ-6k, with their subsets CelebAHQ-250 and FFHQ-600 duplicated 5, 10, 15 and 20 times. Figure 7 depicts the memorization scores of duplicated images. As can be observed, data duplication largely worsens memorization in DMs and our measure is sensitive to these changes. With increasing duplication times, the larger FFHQ-6k dataset presents a stronger resistance to memorization.
For CIFAR-10 and LAION which originally contains duplicates, we observe the effect of deduplication. We deduplicate CIFAR-10 with SSCD embeddings and DBSCAN clustering [14]. This results in 666 identified clusters, only one image of which is retained for training. The memorization is evaluated on the retained 666 images before and after deduplication. On LAION, as compared to SD v1.4, SD v2.1 deduplicates the training set. The memorization is evaluated on the confirmed and suspicous sets. Figure 8 shows the results, with infinite scores highlighted. As can be observed, InvMM clearly exhibits the benefits that deduplication brings.
5.5 Comparison to membership
Figure 9: Performance comparison of membership inference on FFHQ-6k, across different training epochs and duplication times.
A close topic to memorization is membership inference [48, 8, 13], which aims to predict whether a sample is used for training. The common accepted assumption is that member samples are fit “better” than hold-out samples [13], thus statistics like (variant of) training loss or differential training loss [8] are generally used for membership inference. A comparison to a recent membership inference metric SecMI [13] on FFHQ-6k, as shown in Fig. 9, does present a positive correlation between membership and memorization. As factors that exacerbate memorization changes, both SecMI and InvMM shows better performance.
However, there is still a performance gap between membership score and memorization score, which indicates that certain samples from the hold-out set are memorized more than some member samples. This mainly results from the gap between model’s metric (training loss) and human’s metric (perceptual similarity). Membership inference leverages the model’s metric for prediction, while memorization should resort to perceptual metric, e.g., simulated by SSCD. What level of loss leads to image replication depends both on the model and image itself. Low training loss does not necessarily lead to data replication, but would probably distinguish trained data from never accessed. On the contrary, high training loss indicating non-membership may still be enough for replication. In summary, membership does not deterministically lead to high memorization and vice versa. Our measure differs from membership inference metric by focusing on the perceptual level and adaptively handles the specific loss landscape of different samples.
6 Conclusion
In this work, we dive into measuring memorization for diffusion models. Comprehensive experiments demonstrate the ability of the proposed measure to correctly describe the extent to which any image is memorized. In short, an image to a higher extent of memorization is able to be inverted to a more standard and diverse latent distribution. However, our method is also limited by its simple modeling of the sensitive latent noise as multivariate normal. During our research, we observed many clues confirming that the true distribution is complex. Future work might consider more advanced methods, e.g., a generative network, to fit the sensitive noise. Our measure may also support more practical applications, such as online sensitive noise filtering and adversarial training to mitigate harmful memorization.
Supplementary Material
7 Additional technical description
7.1 A unified understanding
Similar to  Eq. 2, the variational lower bound can be further lower bounded w.r.t. a conditional distribution $q_{\bm{\phi}}(\bm{c},\bm{x}_{0})$ with parameters $\bm{\phi}$. For text-to-image models, it is instantiated as a prompt distribution $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$.
Following similar expansion in Eqs. 3 and 4, we obtain the full inversion variational lower bound w.r.t. both condition and latent noise:
$\log p_{\bm{\theta}}(\bm{x}_{0})\geq-l_{de}(\bm{x}_{0};\bm{\phi},\bm{\varphi})-l_{kl}(\bm{x}_{0};\bm{\varphi})-l_{cr}(\bm{x}_{0};\bm{\phi})$ (13)
and
$l_{de}(\bm{x}_{0};\bm{\phi},\bm{\varphi})$ $=-\mathbb{E}_{q_{\bm{\phi}}(\bm{c}|\bm{x}_{0}),q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})}\left[\log p_{\bm{\theta}}(\bm{x}_{0}|\bm{x}_{1},\bm{c})\right.$
$+\sum_{t=2}^{T}$ $\left.D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{x}_{0})\;\|\;p_{\bm{\theta}}(\bm{x}_{t-1}|\bm{x}_{t},\bm{c}))\right]$ (14)
$l_{kl}(\bm{x}_{0};\bm{\varphi})$ $=D_{\mathrm{KL}}(q_{\bm{\varphi}}(\bm{x}_{T}|\bm{x}_{0})\;\|\;p(\bm{x}_{T}))$ (15)
$l_{cr}(\bm{x}_{0};\bm{\phi})$ $=D_{\mathrm{KL}}(q_{\bm{\phi}}(\bm{c}|\bm{x}_{0})\;\|\;p(\bm{c}))$ (16)
where $l_{de}$ indicates the denoising error, $l_{kl}$ is a KL divergence and $l_{cr}$ is a condition regularization term. The three terms can be explained as the following:
  1. 1.
    $l_{de}$ indicates how accurate the pretrained model denoises each $\bm{x}_{t}$ when the added noise is drawn from $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$. If $l_{de}$ is low enough, then for most noises $\bm{\epsilon}_{0}\sim q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ the model presents low denoising error, we can anticipate that the sampling trace starting at $\bm{\epsilon}_{0}\sim q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ will head towards $\bm{x}_{0}$ and finally generate $\bm{x}_{0}$.
  2. 2.
    $l_{kl}$ is a normality regularizer. When $l_{de}$ is optimized to a low level, the noise distribution $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ identifies a sensitive set of noises that will cause the generation of the training image $\bm{x}_{0}$. $p(\bm{x}_{T})$ is the prior distribution, usually set to the standard Gaussian. In a sense, $l_{kl}$ measures the diversity of the model’s generation: When $l_{kl}$ becomes zero, i.e., $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ is standard Gaussian, then low enough $l_{de}$ means the model always generates $\bm{x}_{0}$ and loses generalization.
  3. 3.
    $l_{cr}$ encourages the realistic feasibility of the condition distribution $q_{\bm{\phi}}(\bm{c}|\bm{x}_{0})$. For example, if $p(\bm{c})$ is considered the distribution of natural language, then minimizing $l_{cr}$ indicates that prompt $\bm{c}\sim q_{\bm{\phi}}(\bm{c}|\bm{x}_{0})$ should be grammatically and semantically correct.
In the standard training of diffusion model (DM), $l_{kl}$ is ignored because $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ is set to the standard Gaussian $\mathcal{N}(\mathbf{0},\mathbf{I})$ such that $q_{\bm{\varphi}}(\bm{x}_{T}|\bm{x}_{0})$ approximately equals $\mathcal{N}(\mathbf{0},\mathbf{I})$, $l_{kl}$ approximately equals zero. $q_{\bm{\phi}}(\bm{c}|\bm{x}_{0})$ reduces to several captions coupled with the training image such that $l_{cr}$ is also zero.
Our idea is to measure memorization by relaxation of the noise and prompt distribution so that the denoising error can be optimized low enough to replicate the target image. Based on this, the normality of the worst-case distribution of sensitive latent noise is used as a measure.
7.2 Reparameterization
Sampling from the noise and condition distribution in Eq. 14 is non-differentiable, we indirectly sample them.
When the noise distribution is $q_{\bm{\varphi}}(\bm{\epsilon}_{0}|\bm{x}_{0})$ a multivariate Gaussian $\mathcal{N}(\bm{\mu},\bm{\sigma}^{2})$ with learnable mean $\bm{\mu}$ and diagonal variance $\bm{\sigma}^{2}$,
$\bm{\epsilon}_{0}=\bm{\epsilon}^{\prime}\bm{\sigma}+\bm{\mu},\bm{\epsilon}^{\prime}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (17)
For text-to-image DM, each token $\bm{\omega}_{i}$ (see Sec. 4.4) is reparameterized by
$\tilde{\bm{\omega}}_{i,j}=\frac{\exp{((\log\bm{\pi}_{i,j}+\bm{g}_{i,j})/\tau)}}{\sum_{k=1}^{|\mathcal{V}|}\exp{((\log\bm{\pi}_{i,k}+\bm{g}_{i,k})/\tau)}}$ (18)
where $\{\bm{g}_{i,j}\},i=1...N,j=1...|\mathcal{V}|$ are i.i.d samples drawn from $\mathrm{Gumbel}(0,1)$, $\tau$ is a temperature factor. When $\tau$ approaches 0, the smoothed sample $\tilde{\bm{\omega}}_{i}$ becomes one-hot.
After optimization, we can draw discrete prompt from the learned $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ by:
$\bm{\omega}_{i}=\mathop{\arg\max}_{j}\left[\log\bm{\pi}_{i,j}+\bm{g}_{i,j}\right]$ (19)
8 Additional experiment details and results
8.1 Experiment setting
If not stated otherwise, the hyperparameters follow the default setting listed in Tab. 2. For CelebAHQ and FFHQ, all memorization scores in the plots are evaluated on CelebAHQ-250 and FFHQ-600. All the experiments in this paper are conducted on one NVIDIA A800 GPU.
CIFAR-10 CelebAHQ FFHQ LAION
Learning Rate $\gamma$ 1e-1 1e-1 1e-1 1e-1
Iteration $S$ 2000 2000 2000 2000
Batch Size 32 16 16 16
Cycle $C$ 50 10 10 50
Increment $\delta$ 1e-4 5e-4 5e-4 1e-3
Threshold $\xi$ 1e-3 1e-3 1e-3 1e-3
Threshold $\beta$ 1.0 1.0 1.0 1.0
Sampler DDIM [52] DDIM DDIM DDIM
DDIM Step 200 50 50 50
DDIM $\eta$ 0 0 0 0
Optimizer Adam [27] Adam Adam Adam
SSCD Size 32$\times$32 256$\times$256 256$\times$256 320$\times$320
Table 2: Default hyperparameter settings.
8.2 Detection
Experiment details. On CIFAR-10, the training loss metrics are calculated on the average loss of 16 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. Following van den Burg et al. [54]’s setting, $M_{\mathrm{LOO}}$ is estimated using a 10-fold cross-validation. We only conduct the cross-validation once.
On LAION, the training loss metrics use 32 random Gaussian noise and 50 timesteps uniformly sampled within the range [1, 1000]. We implement Wen et al. [61] following their best performing settings.
Calibration. SSCD similarity on the low-resolution CIFAR-10 yields many false positives (non-replication with high similarity, early stopped) and false negatives (replication with low similarity, not early stopped), although it works well on the other three high-resolution datasets. We perform a manual review on the false samples and correct their early stop timesteps and scores. All the other experimental results are unhandled.
Figure 10: Not invertible examples in SD v1.4. The first column shows the corresponding training images.
Invertibility. More than half of the images in CIFAR-10 are not invertible even if we set $\lambda=0$ all the time. This is probably because the DDPM (34.21 M) trained on CIFAR-10 has limited capacity to memorize every training image. Images in CelebAHQ and FFHQ are all invertible. As a comparison, the LDMs trained on them have larger capacity (314.12 M). There are also some images not invertible in Stable Diffusion. However, setting $\lambda=0$ could invert them almost perfectly. The invertibility of an image is specific to a set of hyperparameter settings. Although it is possible to invert everything in SD, we regard samples not invertible in our setting “insignificant”, as compared to those that are easy to invert. Figure 10 shows three examples that are not invertible in SD v1.4 from the normal subset, as judged by SSCD. It shows that it is quite subjective to judge the similarity between images. Human may regard such cases as replication as the generated images mimic the style of training image or else. From a technical perspective, we leave this problem untouched and follow the decision of SSCD.
8.3 Collate InvMM with nearest neighbor test
Setting CelebAHQ-250 CelebAHQ-2.5k FFHQ-600 FFHQ-6k CelebAHQ FFHQ
Default N/A(0) N/A(0) N/A(0) N/A(0) N/A(0) N/A(0)
Epoch ×$\times 2$ 0.059(9) N/A(0) N/A(0) N/A(0) - -
Epoch ×$\times 3$ 0.326(120) N/A(0) 0.000(1) N/A(0) - -
Epoch ×$\times 4$ 0.567(181) 0.000(3) 0.000(4) N/A(0) - -
Duplicate ×$\times 5$ 0.182(65) - 0.000(2) - -
Duplicate ×$\times 10$ 0.852(237) - 0.192(186) - -
Duplicate ×$\times 15$ 0.937(245) - 0.401(323) - -
Duplicate ×$\times 20$ 0.819(231) - 0.434(375) - -
Table 3: Performance of collating InvMM with nearest neighbor test. Each result refers to IoU(Number of replicated training images).
We provide another quantitative metric to collate InvMM with the nearest-neighbor test: if a training image is replicated by its nearest neighbors in randomly generated samples, then its InvMM should be small among a list of training images, and vice versa. Let $S_{\textrm{nn}}$ be the set of images that expose replication in randomly generated samples, and $S_{\textrm{InvMM}}$ be the same-size set of images with the lowest InvMMs in a list of images. We define the consistence between InvMM and nearest-neighbor test as the Intersection over Union (IoU) between $S_{\textrm{InvMM}}$ and $S_{\textrm{nn}}$:
$\mathrm{IoU}=\frac{|S_{\textrm{InvMM}}\cap S_{\textrm{nn}}|}{|S_{\textrm{InvMM}}\cup S_{\textrm{nn}}|},|S_{\textrm{nn}}|>0$ (20)
Under the setting in Sec. 5.3, InvMM achieves an IoU of 0.817 on CIFAR-10 and 1.0 on LAION. The results on CelebAHQ and FFHQ are summarized in Tab. 3. 10k random samples are used to obtain the results. A random sample of cosine similarity larger than 0.5 with any training sample is considered a replication. InvMM presents consistence with the results of nearest neighbor test, indicating its potential to expose risk of training image leakage, especially when a large number of training images are prone to leakage. The 10k samples is not an adequate sampling of the large latent space (64$\times$64$\times$3), the performance can be refined with a larger scale of evaluation.
8.4 Comparison to membership
Experiment details. Figure 9 is evaluated with LDM trained on FFHQ-6k. FFHQ-600 is used as the member set and CelebAHQ-250 is the hold-out set. The statistic variant of SecMI is implemented following their official settings, with $t_{\mathrm{SEC}}=100$ and $k=10$.
Figure 11: Comparison of replication loss for members and hold-out samples.
Figure 12: Top-2 false positives with lowest InvMM from the hold-out set and top-2 false negatives with highest InvMM from the member set. The results come from LDM trained on FFHQ-6k for 1784 epochs.
Results. Figure 11 provides quantitative evidence that different samples could be replicated with different levels of training loss. The replication loss is calculated over the inverted latent noise distribution. It shows that although holdout samples have a larger training loss over the $\mathcal{N}(0,I)$, supporting effective membership inference, it is still possible to replicate some of them by reducing their training loss to a level higher than some of the members.
We visualize the top-2 false positives and false negatives in Fig. 12. The results show that hold-out samples are also replicated near identically.
The result further highlights the difference between membership and memorization: training loss does not completely determine data replication.
9 Ablation study on prompt inversion
This section elaborates the influence of several factors for prompt inversion in SD, including the temperature $\tau$ in Gumbel-Softmax, the Classifier-Free Guidance (CFG) [21] scale $\gamma$ and the advantage of predicting the image rather than noise for inversion (Eq. 12). For this goal, the noise distribution is temporarily fixed to the standard Gaussian. Experiment results will show that heavily memorized images can also be uncovered in this setting.
Figure 13: Examples in confirmed, suspicious and normal subsets from top to down. In each block, the right two columns show generated images using their training captions but different initial noises.
9.1 Dataset
We use the aforementioned three LAION subsets for evaluation. Figure 13 presents examples from the three subsets, together with images generated using their training captions.
9.2 Experiment setup
We utilize SD v1.4 for evaluation, which uses CLIP [39] to encode input prompts and generate high-resolution $512\times 512$ images. The text encoders by default accepts a maximum length of 77 tokens, in which the first and last tokens are padded tokens indicating the start and end of a prompt. The rest 75 tokens are all optimized in our experiments. During optimization, the parameters $\bm{\theta}$ of diffusion models are fixed. We optimize for 500 iterations with a constant learning rate of 0.2. $\bm{\phi}$ is initialized to $\mathbf{0}$ at the begining of optimization.
9.3 Influence of temperature
Hard prompt inversion to exactly reconstruct certain images is a challenging problem as it requires to search over a large and discrete space consisting of tens of thousands of tokens (49408 in CLIP). We have found that the convergence of inversion relies on appropriate choice of the temperature $\tau$ in Gumbel-Softmax smoothing. With $\tau$ approaching 0, $\tilde{\bm{\omega}}$ drawn from the prompt distribution approaches one-hot and accurately matches a token, while it is difficult to optimize through gradient decent. Larger $\tau$ provides a smoother landscape of the target loss function and thus is easier to optimize. However, the smoothed $\tilde{\bm{\omega}}$ cannot directly correspond to some tokens. Discretizing them anyway (Eq. 19) might not preserve the same effectiveness as the smooth counterparts.
Figure 14: The distribution of denoising error of the confirmed, suspicious and normal subsets. ”dct” means plus discretization.
Denoising error. We first analyze the denoising error of the inverted prompt distribution when optimized with different temperatures. We consider 4 settings for the temperature to be either 0.1, 0.5, 1 or 2, as well as whether to discretize the smoothed tokens $\tilde{\bm{\omega}}$ to $\bm{\omega}$. After optimization, random prompts and noises can be drawn from $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ and $\mathcal{N}(\mathbf{0},\mathbf{I})$. For each image, we randomly sample 10 prompts and 10 noises for each sampled prompt, resulting in 100 prompt-noise pairs. The denoising error is estimated using 50 timesteps uniformly sampled within the range $[1,1000]$ and averaged over the 100 prompt-noise pairs.
The results can be seen in  Fig. 14. For any type of image from different groups, higher temperatures lead to lower denoising errors on average, indicating a more adequate optimization. However, meanwhile, plus token discretization worsens the effectiveness.
(a) Confirmed Examples
(b) Suspicious Examples
(c) Normal Examples
Figure 15: Training denoising error of examples from each group under different temperatures, smoothed via exponential moving average with a momentum of 0.99. ”dct” means plus discretization. The black dashed line is the baseline error calculated using their training captions over 1600 randomly sampled Gaussian noises.
Figure 16: Generation results of different temperatures. The first column shows the corresponding training images.
Convergence.  Figure 15 shows the denoising errors at each optimization step of the 6 example images in  Fig. 13. For the assessment of convergence, we draw a baseline denoising error calculated using the training caption of each image. As can be observed, large temperatures induce better convergence and the difference between $\tilde{\bm{\omega}}$ and $\bm{\omega}$ becomes prominent. Figure 16 illustrates the generation results using prompts sampled from the learned distribution, with a CFG scale of 7. As can be seen, with $\tau=2$, the inverted prompts are able to replicate the training images for the 4 examples from the confirmed and suspicious groups, the two from the suspicious group are newly found through our analysis. However, $\tau=2$ plus discretization produces completely irrelevant images. Lower temperatures 0.5 and 1.0 present more consistent generation between smooth and discrete prompts, while they only produce similar images to the training ones, showing analogous content, color, etc. The smallest $\tau=0.1$ fails to capture the main content of the training images but remains the best consistency for discretization.
Prompt distribution. Figure 17 depicts the density distribution of the entropy $-\sum_{j=1}^{|\mathcal{V}|}\bm{\pi}_{i,j}\log\bm{\pi}_{i,j}$ of the learned prompt categorical distributions. When $\tau=2$, most tokens follow a high entropy distribution, which means that they are well smoothed and take an interpolation of hard tokens. In contrast, smaller temperatures produce more sharp distributions, while less effective as large temperature for inversion.
Figure 17: Entropy density distribution of the prompt categorical distribution.
Conclusion. For the goal of effective analysis, we adopt a compromise setting with the temperature $\tau$ of 2.0 and without discretization, to reach adequate optimization. Although this violates the goal of inverting realistic prompts, it is reasonable and enough for developers to analyze the vulnerability of their models. Note that it still offers a certain level of restriction to the learned soft prompts, as the Gumbel-Softmax approximation together with a linear combination of pretrained token embeddings bound the smoothed tokens $\tilde{\bm{\omega}}$ in the convex hull of the pretrained tokens.
9.4 Alternative prompt tuning methods
Figure 18: Inversion results of AutoPrompt and PEZ. Within each block, left: optimize 75 tokens, right: optimize tokens equal to training captions.
We also compare to several existing discrete prompt tuning methods, to see if they work for training image inversion. Figure 16 has shown that GBDA [20] fails ($\tau=1.0$ with discretization). We further study the effectiveness of AutoPrompt [47] and PEZ [60]. AutoPrompt and PEZ are implemented following their default settings. We evaluate with 4 heavily memorized images (almost 100% sensitivity on their captions). In addition to optimizing all the 75 tokens, we also set the number of optimizable tokens the same as the corresponding training captions (shorter), to reduce the difficulty of inversion. Figure 18 shows the inversion results. AutoPrompt and PEZ both fail to reconstruct the original training images.
Conclusion. Existing discrete prompt tuning methods cannot effectively work to eidetically reconstruct training images. We temporally resort to relaxed prompt tuning.
9.5 CFG Scale
In additional, we sweep the CFG scale $\gamma$ from 0 to 7 with interval 1 to study its influence. $\gamma=0$ indicates unconditional generation and $\gamma=1$ indicates conditional generation without penalizing unconditional prediction. The generation results of the examples in Fig. 13 are shown in Fig. 19.
It can be observed that the generated images with $\gamma=0$ are quite random because they only depend on the random initial noises. When $\gamma\geq 1$, for heavily memorized images in the confirmed and suspicious group, the generation results progressively converge to the training images. At times the generated images with small CFG scale only resemble the training images but are not eidetic, e.g., the 2rd to 4th rows. Nonetheless, we also discovered perfect replication for these examples with a small $\gamma=1$ generated using other different initial noises. This indicates that the extent to which different training images are memorized varies, and, moreover, a relatively low-level training time memorization ($\gamma=1$) can be amplified by sampling-time options such as larger $\gamma$. Given that we optimize the prompts w.r.t. the conditional model ($\gamma=1$), it demonstrates that training data leakage roots in the conditional model.
In addition, a gradual sharpening can be observed in the generated images as the guidance scale increases. As we optimize w.r.t. the conditional model, i.e., $\gamma=1$, it is of enough denosing accuracy to generate an training image with relatively lower scales. Enlarging the conditional scale, however, results in excessive alignment with the input prompt. In contrast, for the images in the normal group (see the last two rows of Fig. 19), as the inverted prompt distribution cannot fully capture its complete content, generation with $\gamma=1$ is somewhat fuzzy. It thus benefits from an increase of $\gamma$ for higher quality.
Conclusion. Training data memorization can be amplified by CFG scale. As we consider the worst-case memorization in this paper, we count in the replication caused by any CFG scale from 1 to 7.
Figure 19: Generation results of different classifier-free guidance scales. The first column shows the corresponding training images.
9.6 Optimization Objective
As we adopt the modified $\bm{x}_{0}$-prediction objective different from the original $\bm{\epsilon}_{0}$-prediction objective of the diffusion models used in our experiments, we verify the effectiveness of $\bm{x}_{0}$-prediction over $\bm{\epsilon}_{0}$-prediction for inversion. We evaluate using the images in the confirmed set to determine if the $\bm{\epsilon}_{0}$-prediction can successfully replicate them. Figure 20 shows the inversion results of $\bm{\epsilon}_{0}$-prediction. Inverison with $\bm{\epsilon}_{0}$-prediction is much unstable compared to $\bm{x}_{0}$-prediction, which demonstrates the importance of reweighting denoising error at different timesteps. More specifically, the later timesteps at training time (ealier at sampling time) tend to shape the large scale image features [22], e.g., shape, object. Therefore, it would be beneficial to upweight the later timesteps by $\bm{x}_{0}$-prediction to more accurately guide diffusion models to generate the corresponding training images.
Conclusion. Although not aligning with the original training objective, $\bm{x}_{0}$-prediction is more stable for inversion.
Figure 20: Inversion results of $\bm{\epsilon}_{0}$-prediction. The first column shows the corresponding training images.
9.7 Adaptive algorithm
Figure 21: Comparison of inversion denoising error under different settings on SD v1.4.
A comparison of the (1) training error (using training captions), (2) inversion error with only prompt distribution learned, (3) inversion error with both prompt and noise distributions learned, fixing $\lambda=1$ and (4) inversion error with both prompt and noise distributions learned, dynamically adjusting $\lambda$ by Algorithm 1, is shown in  Fig. 21. Compared to the training error, (2) only reduces that of heavily memorized images for which the input prompts plays a crucial role. (3) further reduces the denoising error but cannot work for all samples. Algorithm 1 can successfully reduce the denoising error of any training samples by adaptively adjusting the weight of normality regularization.
9.8 Full results
Based on the previous ablation study, we investigate memorization with the appropriate choice of each factor. In addition to SD v1.4, we include SD v2.1 for comparison. When not otherwise specified, experiment settings are the same as ablation study.
Figure 22: Number of training data replication in the confirmed group, sorted in descending order.
We first conduct a complete check of generation results. For each training image, 10 random prompts are sampled from $q_{\bm{\phi}}(\bm{\omega}|\bm{x}_{0})$ and for each sampled prompt, 100 Gaussian noises are sampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$. With an array of CFG scales, each prompt results in 700 randomly generated images. We manually count the number of generated images that present an obvious replication of the corresponding training image and report the maximum number of replications among the 10 prompts.
Figure 22 shows the results of the confirmed set for SD v1.4. There are 61 images being 100% replicated and the rest 15 images are replicated at least 5 times and at most 96 times. In the suspicious set, there are 3 newly discovered images that are replicated 53, 100 and 87 times. This demonstrates the advantage of our inversion-based analysis method, because some images cannot be directly generated by their training captions but do present a relatively large extent of memorization, compared to images in the normal set, where we found no replication.
Figure 23: Inversion results on SD v2.1 (Fixed noise distribution). The first column shows the corresponding training images.
For SD v2.1, we have not observed any replication but very similar generations, shown in Fig. 23. This is probably because SD v2.1 deduplicated the training set, and thus the true likelihood of previous duplicate samples in the training set is significantly reduced.
The counts demonstrate that training images are memorized to varied extent: For some images, training data leakage is certain to happen, for some images, it happens by chance and for the other images, it hardly happens. We also examine the images that present partial replication in our experiments by generating using their training captions and find that indeed, they are not memorized as severely as those that are completely memorized. It thus motivates us to find a measure to tell the difference.
10 More generation results
Figures 24, 25, 26, 27, 28, 29, 30 and 31 show more generated images using inverted noise vectors (and prompts).
Figure 24: Random samples of SD v1.4 inversion on the confirmed subset. The first column shows the corresponding training images.
Figure 25: Random samples of SD v1.4 inversion on the suspicous subset. The first column shows the corresponding training images.
Figure 26: Random samples of SD v1.4 inversion on the normal subset. The first column shows the corresponding training images.
Figure 27: Random samples of SD v2.1 inversion on the confirmed subset. The first column shows the corresponding training images.
Figure 28: Random samples of SD v2.1 inversion on the suspicous subset. The first column shows the corresponding training images.
Figure 29: Random samples of SD v2.1 inversion on the normal subset. The first column shows the corresponding training images.
Figure 30: Random samples of LDM inversion on CelebAHQ. The first column shows the corresponding training images.
Figure 31: Random samples of LDM inversion on FFHQ. The first column shows the corresponding training images.